Finished repairing the CDO this weekend. Due to that repair the STS, Plasma therm and BMR low temp PECVD are back online.
Finished repairs on Hitachi SEM. This system is back online.
We worked on the IM milling, identified a vacuum leak and also replaced the neturalization filament. We are still experiencing problems with the emmision current. We will be addressing that.
Experiencing a problem with the critical point dryer. It is leaking at high pressure. We plan to replace the two retaining rings on the chamber.
We were informed Wednesday that the flood exposure in the INRF UV lamp was not functioning. We recommend if you have a need at this time you utilize the flood exposure in the BiON.
The Thermco 4″ Anneal furnace was contaminated with debris. Furnace tube has been cleaned and is back online.
Users please use the blog to find out status of lab equipment. As equipment comes back online there are two areas to review for status, our website or the blog. We will updating the blog as equipment is back online real time.
Thank you for your patiences.